OpenStudy (anonymous):

solve for x: 2x-3/3 + 5-x/2 = 3

OpenStudy (anonymous):

brake it down into small equations Best way for me to figure it out.

OpenStudy (solomonzelman):

2x-3 and 5-x are the denominators, right? I mean like this, \(\Huge\color{blue}{ \sf \frac{2x-3}{3}+ \frac{5-x}{2} =3}\) right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{(2x-3)}{3} + \frac{(5-x)}{2}=3 }\) \(\Huge\color{blue}{ \sf \frac{(2x-3) \color{red} { \times 2 } }{3 \color{red} { \times 2 } } + \frac{(5-x) \color{red} { \times 3 } }{2 \color{red} { \times 3 } }=3 }\) \(\Huge\color{blue}{ \sf \frac{4x-6}{6} + \frac{15-3x}{6}=3 }\) \(\Huge\color{blue}{ \sf \frac{4x-6}{6} + \frac{15-3x}{6}=\frac{18}{6} }\) what would your next step be ?

OpenStudy (anonymous):

dont know